Circular No. 7847

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bo IN NGC 3190 The Central Bureau has received independent reports of the discovery

of a supernova in NGC 3190 by Paulo Cacella, Brasília, Brazil, on a CCD image taken on Mar. 9.08 UT with a 0.25-m reflector (at mag 16–17), and by Yoji Hirose, Chigasaki, Kanagawa-ken, Japan, on seven unfiltered CCD frames taken ∼ Mar. 9.505 with a 0.25-cm reflector (mag 15.5; communicated by S. Nakano, Sumoto, Japan). Nakano also forwards the following precise position, measured by Y. Kushida from an unfiltered CCD image taken by R. Kushida at Yatsugatake South Base Observatory on Mar. 9.604 (with SN 2002bo at mag 14.6) with a 0.40-m reflector: α = 10^{h}18^{m}06^{s}.51, δ = +21^{o}49′41′′.7 (equinox 2000.0), or 11′′.6 east and 14′′.2 south of the galaxy’s center; nothing is visible at this location on the Digitized Sky Survey or on Kushida’s numerous images of this galaxy dating back to 1995, including 2002 Feb. 14 (limiting mag 18). Hirose reported the new object to be 17′′ east and 23′′ south of the center of NGC 3190, and Cacella provided position end figures 06^{s}.45, 40′′.2. W. D. Li, University of California at Berkeley, reports that SN 2002bo is present on a CCD image obtained with the Katzman Automatic Imaging Telescope on Mar. 9.4 at mag ∼ 16; he provides position end figures 06^{s}.52, 41′′.2. A KAIT image taken on Mar. 2.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002bp IN UGC 6332 T. Puckett and A. Langoussis, Mountain Town, GA, report the dis-

covery of an apparent supernova (mag 18.6) on an unfiltered CCD frame (limiting mag 20.3) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Mar. 8.16 UT. The new object is located at α = 11^{h}19^{m}18^{s}.20, δ = +20^{o}48′23′′.1 (equinox 2000.0), which is 21′′.3 east and 26′′.0 south of the center of UGC 6332. SN 2002bp was confirmed on CCD frames taken on Mar. 9.14, but it is not present on images taken by Puckett on 2001 Mar. 19, 28, and May 15 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1995 Mar. 26 (limiting mag ∼ 21.0), 1992 May 1 (limiting mag ∼ 21.0), and 1955 Mar. 22 (limiting mag ∼ 20.0).

NOVA IN THE LARGE MAGELLANIC CLOUD 2002 Photoelectric photometry by P. Kilmartin and A. C. Gilmore, Mt. John

Observatory: Mar. 4.413 UT, V = 11.16, U − B = −0.39, B − V = +0.42, V − R = +0.32, V − I = +0.69; 5.403, 11.29, −0.52, +0.35, +0.37, +0.79.

2002 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green